# Evaluating the Bias in LLMs for Surveying Opinion and Decision Making in Healthcare

### v1

@@ Abstract
Generative agents have been increasingly used to simulate human behaviour in silico, driven by large language models (LLMs). These simulacra serve as sandboxes for studying human behaviour without compromising privacy or safety. However, it remains unclear whether such agents can truly represent real individuals. This work compares survey data from the Understanding America Study (UAS) on healthcare decision-making with simulated responses from generative agents. Using demographic-based prompt engineering, we create digital twins of survey respondents and analyse how well different LLMs reproduce real-world behaviours. Our findings show that some LLMs fail to reflect realistic decision-making, such as predicting universal vaccine acceptance. However, Llama 3 captures variations across race and Income more accurately but also introduces biases not present in the UAS data. This study highlights the potential of generative agents for behavioural research while underscoring the risks of bias from both LLMs and prompting strategies.
@@
Taylor Anderson3 tander6@gmu.edu                        Matthew Scotch4 matthew.scotch@asu.edu                        David Heslop1 d.heslop@unsw.edu.au
@@ 2 Method
We evaluate vaccination decisions by testing models across four historical pandemic contexts from 2020. Each model is presented with a standardised decision-making prompt, incorporating demographic details and situational factors. Model predictions are then compared to UAS survey data to assess alignment with real-world trends across different pandemic phases. To analyse biases, we examine disparities in LLM-generated decisions within each demographic category (e.g., variations in vaccine hesitancy across racial or Income groups). To understand potential biases, we analyse both internal inconsistencies across demographic groups and compare model outputs to real-world survey data (UAS) where available. This helps identify whether LLMs reflect known behavioural differences or exaggerate group-level patterns. This approach assesses LLMs’ reliability in healthcare decisions and highlights potential biases from demographic variations in responses.
@@ 2.2.1 Experimental Setup
We assess each LLM to see how contextual variations influence decision-making. We also analyse bias to identify disparities in LLM responses regarding vaccine hesitancy and whether demographic details mitigate biases. Each model generates 11.5k samples across all demographic profiles and four pandemic phases. We apply majority voting over three generations per prompt to reduce response variance. In designing the prompt, we experimented with different prompt formats, including persona-only, location-based, situation-based, and few-shot prompts. However, few-shot prompts often made the model repeat or exaggerate specific examples, especially when sensitive attributes like race or Income were included. To reduce this risk, we chose a zero-shot prompt with contextual information, which gave more consistent and balanced results.
@@ 3.1 Comparison of LLM predictions with UAS survey data
Our study examines vaccination intentions using four LLMs, prompting them with pandemic-phase-specific contexts and demographic profiles. The structured prompts cover four COVID-19 phases in the U.S. (Jan–Dec 2020), allowing us to assess how LLMs simulate decision-making trends compared to UAS survey data (Figure 2). Vaccine hesitancy is defined as the percentage of responses where the model (or survey participant) answered that they would not get vaccinated.
@@ 3.2 Bias analysis in LLM-generated decisions
We define bias in this context not as the presence of group-level differences, but as distortions or misalignment from real-world trends. Our goal is to examine whether LLM-generated disparities reflect known behavioural patterns or diverge in ways that may misrepresent specific groups. Our results reveal significant disparities in vaccine decisions generated by LLMs across income, education, and race, as shown in Table 1. Gemma2 and Ministral exhibit the most pronounced disparities, with low DIR values for income (0.236 and 0.404) and education (0.061 and 0.349). Their high JSD scores also indicate substantial divergence across groups. Prior work has shown that income and education correlate with vaccine hesitancy Aw et al. (2021); Allen et al. (2021), but the gap in model outputs suggests a disconnect between LLM behaviour and real-world trends.
Galactica produces the most balanced predictions across demographic groups, while Llama3 performs moderately but tends to underestimate hesitancy overall. Among the demographic dimensions, we focus on race in Figure 3 due to the distinct trends observed across models. The rightmost bars in the figure represent UAS survey data, which shows higher vaccine hesitancy among Black respondents compared to White and Asian groups. Most LLMs, however, flatten this distribution. For instance, Llama3 shows uniformly low hesitancy across races, and Galactica and Ministral generate nearly equal predictions for all groups. Gemma2 reflects the real-world ranking more closely and predicts the highest hesitancy for Black individuals—over 70%—but still reduces the contrast compared to the UAS survey.

### v2

@@ Abstract
Generative agents have been increasingly used to simulate human behaviour in silico, driven by large language models (LLMs). These simulacra serve as sandboxes for studying human behaviour without compromising privacy or safety. However, it remains unclear whether such agents can truly represent real individuals. This work compares survey data from the Understanding America Study (UAS) on healthcare decision-making with simulated responses from generative agents. Using demographic-based prompt engineering, we create digital twins of survey respondents and analyse how well different LLMs reproduce real-world behaviours. Our findings show that some LLMs fail to reflect realistic decision-making, such as predicting universal vaccine acceptance. However, Llama 3 captures variations across race and Income more accurately but also introduces biases not present in the UAS data. This study highlights the potential of generative agents for behavioural research while underscoring the risks of bias from both LLMS and the prompting strategy.
@@
Taylor Anderson3 tander6@gmu.edu                        C. Raina MacIntyre1 r.macintyre @unsw.edu.au                        Matthew Scotch4 matthew.scotch@asu.edu                        David Heslop1 d.heslop @unsw.edu.au
@@ 2 Method
We evaluate vaccination decisions by testing models across four historical pandemic contexts from 2020. Each model is presented with a standardised decision-making prompt, incorporating demographic details and situational factors. Model predictions are then compared to UAS survey data to assess alignment with real-world trends across different pandemic phases. To analyse biases, we examine disparities in LLM-generated decisions within each demographic category (e.g., variations in vaccine acceptance across racial or Income groups). To understand potential biases, we analyse both internal inconsistencies across demographic groups and compare model outputs to real-world survey data (UAS) where available. This helps identify whether LLMs reflect known behavioural differences or exaggerate group-level patterns. This approach assesses LLMs’ reliability in healthcare decisions and highlights potential biases from demographic variations in responses.
@@ 2.2.1 Experimental Setup
We assess each LLM to see how contextual variations influence decision-making. We also analyse bias to identify disparities in LLM responses regarding vaccine acceptance and whether demographic details mitigate biases. Each model generates 11.5k samples across all demographic profiles and four pandemic phases. We apply majority voting over three generations per prompt to reduce response variance. In designing the prompt, we experimented with different prompt formats, including persona-only, location-based, situation-based, and few-shot prompts. However, few-shot prompts often made the model repeat or exaggerate specific examples, especially when sensitive attributes like race or Income were included. To reduce this risk, we chose a zero-shot prompt with contextual information, which gave more consistent and balanced results.
@@ 3.1 Comparison of LLM predictions with UAS survey data
Our study examines vaccination intentions using four LLMs, prompting them with pandemic-phase-specific contexts and demographic profiles. The structured prompts cover four COVID-19 phases in the U.S. (Jan–Dec 2020), allowing us to assess how LLMs simulate decision-making trends compared to UAS survey data (Figure 2). vaccine acceptance is defined as the percentage of responses where the model (or survey participant) answered that they would not get vaccinated.
@@ 3.2 Bias analysis in LLM-generated decisions
We define bias in this context not as the presence of group-level differences, but as distortions or misalignment from real-world trends. Our goal is to examine whether LLM-generated disparities reflect known behavioural patterns or diverge in ways that may misrepresent specific groups. Our results reveal significant disparities in vaccine decisions generated by LLMs across income, education, and race, as shown in Table 1. Gemma2 and Ministral exhibit the most pronounced disparities, with low DIR values for income (0.236 and 0.404) and education (0.061 and 0.349). Their high JSD scores also indicate substantial divergence across groups. Prior work has shown that income and education correlate with vaccine acceptance Aw et al. (2021); Allen et al. (2021), but the gap in model outputs suggests a disconnect between LLM behaviour and real-world trends.
Galactica produces the most balanced predictions across demographic groups, while Llama3 performs moderately but tends to underestimate hesitancy overall. Among the demographic dimensions, we focus on race in Figure 3 due to the distinct trends observed across models. The rightmost bars in the figure represent UAS survey data, which shows higher vaccine acceptance among Black respondents compared to White and Asian groups. Most LLMs, however, flatten this distribution. For instance, Llama3 shows uniformly low hesitancy across races, and Galactica and Ministral generate nearly equal predictions for all groups. Gemma2 reflects the real-world ranking more closely and predicts the highest hesitancy for Black individuals—over 70%—but still reduces the contrast compared to the UAS survey.